OpenStudy (anonymous):

Identify the sequence graphed below and the average rate of change from n = 1 to n = 3.

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

@dan815 @ParthKohli @jdoe0001 @jinxhead20

OpenStudy (anonymous):

Can you see that for every next term original (x,y) next term (x+1,y/2) ? Do you understand what I am saying?

OpenStudy (anonymous):

YEs

OpenStudy (anonymous):

(x,y) --> (x+1,y/2) That is the pattern that I see

OpenStudy (anonymous):

Not one of the choices is it?

OpenStudy (anonymous):

n=1 and n=3 are you talking about the 1st and 3rd term?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Is the answer A?

OpenStudy (anonymous):

I really don't know what they would want from you when they say "identify" the sequence I also thought the same thing, A

OpenStudy (anonymous):

I think it IS A

OpenStudy (anonymous):

Can I ask some more questions?

OpenStudy (anonymous):

Ramona is playing a board game and rolls two number cubes. Let A = {the sum of the number cubes is odd} and let B = {the sum of the number cubes is divisible by 5}. List the outcomes in A ∩ B

OpenStudy (anonymous):

I'm not sure about this one...

OpenStudy (anonymous):

I got a lot going on so let me just tag @Destinymasha

OpenStudy (anonymous):

Good luck...

OpenStudy (zzr0ck3r):

the sequence is \(8(\frac{1}{2})^{n-2}\) do you see why?

OpenStudy (anonymous):

I know why it was because I plugged it in to the equation: Ramona is playing a board game and rolls two number cubes. Let A = {the sum of the number cubes is odd} and let B = {the sum of the number cubes is divisible by 5}. List the outcomes in A ∩ B

OpenStudy (zzr0ck3r):

and from from 1 to 3 the avg rate is \(\frac{a_3-a_1}{3-1}=\frac{4-16}{2}=-6\)

OpenStudy (anonymous):

Ramona is playing a board game and rolls two number cubes. Let A = {the sum of the number cubes is odd} and let B = {the sum of the number cubes is divisible by 5}. List the outcomes in A ∩ B

OpenStudy (anonymous):

This was the question I asked you

OpenStudy (zzr0ck3r):

good luck with that attitude, it will work out